Find the sum of: 
137 and – 354 ?????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

137 and -354
(137) + (-354)
= (317) + (-137) + (-217)
= 0 + (-217) [(317) + (-317) = 0]
= (-217)

Thanks